# About possible measures in Quantum Gravity

### v2

@@ Abstract
Possible measures for Quantum Gravity are considered. Choices that are invariant under diffeomorphisms are analyzed, but the possibility of employing non invariant measures is also taken into account. The last possibility may be accepted if the anomaly in the measure is compensated by counter term redefinitions of the model under analysis. Particular attention is paid to some concrete examples of non covariant looking measures, which may be useful for generalizing the Veltmann identities when quantizing gravity around curved space times. The results are specified for the Stelle gravity model [1]-[2], which is known to be renormalizable in flat space, although not known to be so in curved ones.
@@ 1. Introduction
If one is able to find a perfectly invariant measure under the fundamental symmetries of the model, that is $A(f)=1$, then this is a useful result. If not, it does not necessarily implies that the theory must be abandoned. Furthermore, there may be several different invariant measures and are equally valid from the theoretical point of view. The allowance of these measures depends on the model, since the anomaly term has to match the counter terms of the corresponding lagrangian.
@@
$d\mu_{c}(\phi,g_{\mu\nu})=\prod d\widetilde{\phi},\qquad\widetilde{\phi}=g^{\frac{1}{3}}\phi.$
@@
with $a$, $b$, $c$, and $d$ are arbitrary real numbers. This arbitrary choice implies that $\widetilde{\phi}$ and $\widetilde{\pi}$ are not necessarily canonically conjugated variables. For a large set of these real numbers, the measure is anomalous under general coordinate transformations. Despite this anomaly, the outcome of reference [13] is that the anomalies correspond to different choices of $a$, $b$, $c$, and $d$ are related all by redefinitions of the counterterms that can be added to the action. This indicates equivalence between the different choices. This means that it is possible, for an scalar field $\phi$, to ensure invariance under coordinate transformations and simultaneously, to ensure unitarity, up to a counter term redefinition.
@@
The intention of the present work is to study apparently non covariant measures, such as the ones in [44]-[45], in connection with the cancellation of volume divergences $\delta^{4}(0)$.
@@ 3.1 A preliminary higher derivative model
which represents $\dot{\phi}_{a}$ up to functions solely on $\phi_{b}$. This reduces the order of the corresponding system. The type of transformations just described in (3.13), though limited, will be useful for the gravity scenario to be considered later on. The last relation can be inverted
@@
After simplifying some terms and performing the Gaussian integration on $\Pi^{a}$ and $P^{a}$ by use of the Gaussian formula
it is found
In other words
Cancellations of the divergent terms proportional to $\delta^{4}(0)$.
This measure can be expressed in exponential form as follows
$d\mu_{L}(\phi_{a})=\prod_{a,x}d\phi^{a}(x)\prod_{x,a}\exp\bigg\{\frac{1}{2}\log(\det(-A^{ab}(\phi)))\bigg\}\exp\bigg\{\frac{1}{2}\log(\det(D^{ab}(\phi)))\bigg\}$
$\prod_{a,x}d\phi^{a}(x)\prod_{a}\exp\bigg\{\frac{1}{2}\sum_{x}\log(\det(-A^{ab}(\phi)))\bigg\}\exp\bigg\{\frac{1}{2}\sum_{x}\log(\det(D^{ab}(\phi)))\bigg\}$
$\lim_{\epsilon\to 0}\prod_{a,x}d\phi^{a}(x)\prod_{a}\exp\bigg\{\frac{1}{2}\frac{1}{\epsilon^{4}}\sum_{x}\log(\det(-A^{ab}(\phi)))\epsilon^{4}\bigg\}\exp\bigg\{\frac{1}{2}\frac{1}{\epsilon^{4}}\sum_{x}\log(\det(D^{ab}(\phi)))\epsilon^{4}\bigg\}$
$\prod_{a,x}d\phi^{a}(x)\prod_{a}\exp\bigg\{\frac{1}{2}\delta^{4}(0)\int\log(\det(-A^{ab}(\phi)))d^{4}x\bigg\}\exp\bigg\{\frac{1}{2}\delta^{4}(0)\int\log(\det(D^{ab}(\phi)))d^{4}x\bigg\}.$ (3.15)
The sign $|_{0}$ indicates that the functional derivatives are evaluated on the classical solutions. The second term vanishes, since the classical equations of motion are satisfied. The last, for the theories under consideration, can be worked out as
$Z=e^{iS_{0}}Z_{1},$
where $Z_{1}$ is related to the determinant of the inverse of the operator
$\det(O)=\det(U)\det(Z-WU^{-1}V).$
$\text{Tr}\,O=\int d^{4}x\lim_{y\to x}O\,\delta^{4}(x-y),$
$\delta^{4}(x-y)=\sum_{n}\psi_{n}(x)\psi_{n}^{\ast}(y),$
$\text{Tr}\,O=\sum_{n}\lambda_{n},$
3.2 Quadratic gravity
$ds^{2}=-dt^{2}+g_{ij}(x)dx^{i}dx^{j},$
$K_{ij}=\frac{1}{2}\dot{g}_{ij},$ (3.18)
the lagrangian (2.8) takes the following form
3.3 The case for quadratic gravity
Introduce the additional variable
$K_{ij}=\frac{1}{2N}\left(\dot{g}_{ij}-2\nabla_{(i}N_{j)}\right)\,.$ (3.20)
$(E+L)^{2}-2V_{i}V^{i}+\beta\left(2E+R+L\right)^{2}\Big]\,$,
$L=K_{ij}K^{ij}-K^{2}\qquad V_{i}=\nabla^{j}K_{ij}-\nabla_{i}K,$
The time derivative of the metric can be expressed in terms of $K_{ij}$ as
$\dot{g}_{ij}=2NK_{ij}+2\nabla_{(i}N_{j)}.$
$G^{ijkl}=\alpha(g^{ik}g^{jl}+g^{il}g^{jk})+2\upsilon_{4}g^{ij}g^{kl},$
$\dot{K}_{ij}=\frac{N}{\sqrt{g}}\tilde{Q}_{ij}+B_{ij}.$ (3.22)
where
$Q^{ij}=P^{ij}-2\sqrt{g}\left(\alpha R^{ij}-g^{ij}X\right)$
$J_{ijkl}=2K_{ik}K_{jl}-K_{ij}K_{kl}\,,\quad J_{ij}=g^{kl}J_{ijkl}.$
the ghost action is
$S_{g}=\int\sqrt{g}\sqrt{h}[h^{ij}b_{i\alpha}\partial_{j}c^{\alpha}+g^{\alpha\beta}b_{\alpha j}\partial_{\beta}c^{j}]d4x.$
[…]
4. Discussion
In the present note, possible invariant measures for Quantum Gravity were considered. The first type of measure studied are invariant under coordinate transformations, and mimics the properties considered in [47], [48]-[52] based on the earlier work [4]. In this formalism, the fact that the measure is invariant under general coordinate transformations is justified without relying on the gravity model in consideration. The approach is geometric and introduces a natural notion of distance in the space of metric component configurations leading to the invariant measure. These types of measures can be applied to almost any quantum theory of gravity, since its invariance under coordinate transformations can be justified by its purely geometric arguments. In four dimensions, the simplest choice was derived in (2.6). However, in these cases, there is a propagating $\delta^{4}(0)$ term in the action which complicates the renormalization program. Of course, when quantization along flat space is considered, and dimensional regularization is employed, then the Veltman identities set these quantities to zero. In curved spaces, for higher loops, these identities or some alternative covariant method of quantization allowing the avoidance of volume divergences has to be worked out independently.
On the other hand, there exist a series of measures, such as the ones presented in [7], [44]-[45], whose expression strongly depends on the model under study. The invariance of these measures is justified by assuming that the Feynman path integral $Z$ is invariant under different choices of canonical transformations. This is applied in order to determine the measure and its change under general coordinate transformation, by assuming that in all the intermediate steps the action is given by a generic type of lagrangian such as (2.9). All the steps involved are dependent on the lagrangian in consideration, and so does the measure. The result contains factors in powers of $g^{00}$. The argument however may be dubious if there appears an anomaly invalidating the general form (2.9).
Although the factors $g^{00}$ are apparently non covariant and possibly anomalous, they can be useful for the elimination of the divergences proportional to $\delta^{4}(0)$ in the effective action. In the present work a generalization of the measure found in [44] was shown to cancel these unwanted terms $\delta^{4}(0)$, at least in the extremal, for a large but not completely general type of gauges (the ones for which the full gravity plus ghost action takes the form (3.12) with both usual and Grassmann variables). These measures have been source of debate and their use in practical applications was avoided in several references [59]-[61]. While this approach is grounded on concrete calculations and can not be criticized, the authors opinion is that the arguments so far do not constitute an elimination point. There are works which instead, prefer the use of the non flat measure [26]-[30]. In fact, these measures may be the key for generalizing the Veltman identities in a general context, where perturbation theory is formulated over a general curved background. These apparently non covariant measures can be discarded only if it is proved that their anomaly under general coordinate transformations can not be adsorbed with the allowed counterterms of the model. Since the full counterterm structure of Stelle gravity, perturbed around curved backgrounds is not yet established, it is difficult to reject them. In fact, they may provide one of the nice features of dimensional regularization for curved backgrounds namely, the vanishing of the volume $\delta^{4}(0)$ divergencies.
Note that the method [7], when applied to an scalar fields, give rise to an anomalous measure. However, this anomalous measure induce counter term redefinitions and is perfectly acceptable [13]. It would be a nice feature if this was the case in gravity, suggesting the avoidance of the divergent $\delta^{4}(0)$ terms from the very beginning.
It is perhaps worthy to consider general measures depending on powers of $g^{a}$ and $(g^{00})^{b}$, and to analyze the propagation of these factors in specific calculations, instead of throwing them away from the very beginning.
Acknowledgements
The author is supported by CONICET, Argentina.

### v3

@@ Abstract
There exist several different proposals for a measure in Quantum Gravity theories. Although sometimes being labelled as non covariant, the measure derived in [7] for GR has the particularity that, in the extremal, the volume divergences proportional to $\delta^{4}(0)$ cancel out. The analogous for Quadratic Gravity [1]-[2] was considered in [44]-[45]. However, as far as the author’s knows, the issue of volume divergences was not considered for this last measure. The present work fills this gap and presents an analysis showing that, in the extremal, these divergences cancel as well. This is up to some subtleties related to superdeterminants. The possibility of employing non invariant measures may be accepted if the anomaly in the measure is compensated by counter term redefinitions of the model under analysis. This makes difficult to disprove, at the present times, some choices of measures. Quadratic Gravity[1]-[2], is known to be renormalizable in flat space, and there are a finite number of counter terms needed in order to renormalize its effective action. However, around a curved space this is not known, and this complicates considerably the analysis. These issues are reviewed in the text, together with an analysis of covariant measures. In particular, it is shown how these measures [47]-[49] can be found if one condition in [7] is relaxed.
@@ 1. Introduction
If one is able to find a perfectly invariant measure under the fundamental symmetries of the model, that is $A(f)=1$, then this is a useful result. If not, it does not necessarily implies that the theory must be abandoned. Furthermore, there may be several different invariant measures equally valid from the theoretical point of view. The allowance of these measures depends on the model, since the anomaly term has to match the counter terms of the corresponding lagrangian.
@@
$d\mu_{c}(\phi,g_{\mu\nu})=\prod d\widetilde{\phi},\qquad\widetilde{\phi}=g^{\frac{1}{4}}\phi.$
@@
with $a$, $b$, $c$, and $d$ are arbitrary real numbers. This arbitrary choice implies that $\widetilde{\phi}$ and $\widetilde{\pi}$ are not necessarily canonically conjugated variables. For a large set of these real numbers, the measure is anomalous under general coordinate transformations. Despite this anomaly, the outcome of reference [13] is that the anomalies correspond to different choices of $a$, $b$, $c$, and $d$ are related all by redefinitions of the counterterms that can be added to the action. This indicates equivalence between the different choices. This means that it is possible, for an scalar field $\phi$, to ensure invariance under coordinate transformations and simultaneously, to ensure unitarity, up to a counter term redefinitions.
@@
The intention of the present work is to study apparently non covariant measures, such as the ones in [44]-[45], in connection with the cancellation of volume divergences $\delta^{4}(0)$. It is well known that these divergences are cancelled in GR [7], at least at the extremal. However, to the best of the authors knowledge, this analysis has not been done for the measures [44]-[45]. The present work fills this hole, and provides an analysis indicating that this cancellation also holds for Quadratic Gravity. The arguments presented in the text of course, do not imply that this measure is one of the correct ones. But it points out one of the interesting features of dimensional regularization namely, the disappearance of volume divergences. In this sense, the measure found in [44]-[45] is special.
@@ 2.1 A model independent invariant measure for gravity
At this point, no regularization of $\delta^{4}(0)\int\partial_{\mu}\epsilon^{\mu}d^{4}x$ is presented. The reason is that if the terms of this form are cancel each other out before regularization, then it is not needed to specify it. This is precisely one of the conditions to be requested below.
@@ 3.1 A preliminary higher derivative model
which represents $\dot{\phi}_{a}$ up to functions solely on $\phi_{b}$. The lagrangian has to be expressed in terms of these new variables. The Gauss-Ostrogradksy method for dealing with these theories in hamiltonian form consists of introducing canonically conjugated momentum densities to $\phi_{a}$ and $\psi_{a}$ given respectively by
The standard classical Poisson brackets lead to the same equations of motion than in the lagrangian formalism, for these higher order theories [38]-[40]. The resulting hamiltonian is linear in $\pi^{a}$ and the Hamiltonian path integral formulation of the model is divergent. The references [36]-[40] propose to distinguish between the variables which are even under time reversal $TqT^{-1}=q$ and those which are odd $TqT^{-1}=-q$. For the first case one should implement usual quantization. For the second type of variables, known as the Dirac-Pauli type variables [55]-[56], one has to employ a variant quantization.
In the following, the result of the above procedure will be presented. The step by step details are collected in the appendix.
After applying this quantization procedure (see full details in the appendix) the last expression becomes
or explicitly
$\lambda^{a}(\phi_{a}-F_{a}(\phi^{b},\eta^{a}))+\beta^{a}(\eta_{a}-G_{a}(\phi^{b},\eta^{a}))]d^{4}x\bigg\}.$ (3.14)
All the intermediate steps leading to this formula are collected in the appendix. After simplifying some terms and performing the Gaussian integration on $\Pi^{a}$ and $P^{a}$ by use of the Gaussian formula
it is found
In other words
Cancellations of the divergent terms proportional to $\delta^{4}(0)$.
This measure can be expressed in exponential form as follows
$d\mu_{L}(\phi_{a})=\prod_{a,x}d\phi^{a}(x)\prod_{a}\exp\bigg\{\frac{1}{2}\delta^{4}(0)\int\log(\det(-A^{ab}(\phi)))d^{4}x\bigg\}$
$\exp\bigg\{\frac{1}{2}\delta^{4}(0)\int\log(\det(D^{ab}(\phi)))d^{4}x\bigg\}.$ (3.15)
The sign $|_{0}$ indicates that the functional derivatives are evaluated on the classical solutions. The second term vanishes, since the classical equations of motion are satisfied. The last, for the theories under consideration (3.12), can be worked out as
$Z=e^{iS_{0}}Z_{1},$
where $Z_{1}$ is related to the determinant of the inverse of the operator
$\det(O)=\det(U)\det(Z-WU^{-1}V).$
$\text{Tr}\,O=\int d^{4}x\lim_{y\to x}O\,\delta^{4}(x-y),$
$\delta^{4}(x-y)=\sum_{n}\psi_{n}(x)\psi_{n}^{\ast}(y),$
$\text{Tr}\,O=\sum_{n}\lambda_{n},$
4. Quadratic gravity
4.1 Synchronous gauge
$ds^{2}=-dt^{2}+g_{ij}(x)dx^{i}dx^{j},$
$K_{ij}=\frac{1}{2}\dot{g}_{ij},$ (4.18)
the lagrangian (2.8) takes the following form
4.2 Other gauges and superdeterminants
Introduce the additional variable
$K_{ij}=\frac{1}{2N}\left(\dot{g}_{ij}-2\nabla_{(i}N_{j)}\right)\,.$ (4.20)
$(E+L)^{2}-2V_{i}V^{i}+\beta\left(2E+R+L\right)^{2}\Big]\,$,
$L=K_{ij}K^{ij}-K^{2}\qquad V_{i}=\nabla^{j}K_{ij}-\nabla_{i}K,$
The time derivative of the metric can be expressed in terms of $K_{ij}$ as
$\dot{g}_{ij}=2NK_{ij}+2\nabla_{(i}N_{j)}.$
$G^{ijkl}=\alpha(g^{ik}g^{jl}+g^{il}g^{jk})+2\upsilon_{4}g^{ij}g^{kl},$
$\dot{K}_{ij}=\frac{N}{\sqrt{g}}\tilde{Q}_{ij}+B_{ij}.$ (4.22)
where
$Q^{ij}=P^{ij}-2\sqrt{g}\left(\alpha R^{ij}-g^{ij}X\right)$
$J_{ijkl}=2K_{ik}K_{jl}-K_{ij}K_{kl}\,,\quad J_{ij}=g^{kl}J_{ijkl}.$
the ghost action is
$S_{g}=\int\sqrt{g}\sqrt{h}[h^{ij}b_{i\alpha}\partial_{j}c^{\alpha}+g^{\alpha\beta}b_{\alpha j}\partial_{\beta}c^{j}]d4x.$ (4.24)
[…]
5. Discussion
In the present work it was shown that the measure [44]-[45] for Quadratic Gravity has been shown to ensure cancellation of the volume divergences $\delta^{4}(0)$, at least in the extremal. This is up to subtle details related to superdeterminants, discussed below formula (4.24). This imitates the features found in [7] for their measure in GR.
The above result alone of course, does not imply that these measures are the correct ones. They only imply that these measures posses in the extremal a nice property which is usual in dimensional regularization namely, the vanishing of the terms proportional to $\delta^{4}(0)$. This fact is not mandatory, it comes as a theoretical bonus. Note that there exist references such as [71]-[72] which consider the possibility of propagating these divergences and dealing with them with counter terms. This approach is perfectly legitimate, however, it may be hard to apply.
For this reason, other possible invariant measures for Quantum Gravity were considered. The first type of measure studied are invariant under coordinate transformations, and mimics the properties considered in [47], [48]-[52] based on the earlier work [4]. In this formalism, the fact that the measure is invariant under general coordinate transformations is justified without relying on the gravity model under study. The approach is geometric and introduces a natural notion of distance in the space of metric component configurations leading to the invariant measure. These types of measures can be applied to almost any quantum theory of gravity, since its invariance under coordinate transformations can be justified by its purely geometric arguments. In four dimensions, the simplest choice was derived in (2.6). However, in these cases, there is a propagating $\delta^{4}(0)$ term in the action which complicates the renormalization program. Of course, when quantization along flat space is done, and dimensional regularization is employed, then the Veltman identities set these quantities to zero. In curved spaces, for higher loops, these identities or some alternative covariant method of quantization allowing the avoidance of volume divergences has to be worked out independently.
It is important to remark that formally, there is no reason for requesting that the volume divergences cancel out. One may work a renormalization program which takes care of them independently. However, this may be hard.
A crucial difference between the measures considered by the authors [47], [48]-[52] and [4] with the ones found in [7], [44]-[45], is that the last ones strongly depend on the model under study. The invariance of these measures is justified by assuming that the Feynman path integral $Z$ is invariant under different choices of canonical transformations. This is applied in order to determine the measure and its change under general coordinate transformation, by assuming that in all the intermediate steps the action is given by a generic type of lagrangian such as (2.9). All the steps involved are dependent on the lagrangian in consideration, and so does the measure. The result contains factors in powers of $g^{00}$. The argument however may be dubious if there appears an anomaly invalidating the general form (2.9).
These measures have been source of debate and their use in practical applications was avoided in several references [59]-[61]. While this approach is grounded on concrete calculations and can not be criticized, the authors opinion is that the arguments so far do not constitute an elimination point. There are works which instead, prefer the use of the non flat measure [26]-[30]. These apparently non covariant measures can be discarded only if it is proved that their anomaly under general coordinate transformations can not be adsorbed with the allowed counterterms of the model. Since the full counterterm structure of Stelle gravity, perturbed around curved backgrounds is not yet established, it is difficult to reject them.
Note that the method [7], when applied to an scalar fields, give rise to an anomalous measure. However, this anomalous measure induce counter term redefinitions and is perfectly acceptable [13]. It would be a nice feature if this was the case in gravity, suggesting the avoidance of the divergent $\delta^{4}(0)$ terms from the very beginning. The same follows for the measures of [44]-[45]. It is perhaps worthy to consider general measures depending on powers of $g^{a}$ and $(g^{00})^{b}$, and to analyze the propagation of these factors in specific calculations, instead of throwing them away from the very beginning.
Acknowledgements
The author is supported by CONICET, Argentina and by the Grant PICT 2020-02181.
Appendix A Some explicit calculations
In this appendix, the explicit pass from the lagrangian (3.12) to the expression (3.14) will be derived. From this lagrangian and (3.13) it is seen that
@@ Appendix A Some explicit calculations
This is exactly (3.14).